# Global well-posedness of the elastic-viscous-plastic sea-ice model with the inviscid Voigt-regularisation

### 06 Mar 2026 v2

@@ 3.3 Well-posedness of the intermediate system (13)
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (76)
$\mathcal{D}_{i}$ $=\lvert D(u_{i})\rvert$. (77)
Subsequently, by proceeding entirely analogously to section 3.3 (the only difference being that we use Lemma 2.3 in the case $\epsilon=0$ rather than $\epsilon>0$), we can obtain the analogous version of estimate (68), which is given by
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (78)
@@
We remark that the proof of the convergence of the solutions in the viscosity cutoff limit $\epsilon\rightarrow 0$, is in some sense related to the approach in Ref. [3], in which the local well-posedness of a model for non-Newtonian incompressible flows with power-law stress tensor was established uniformly in the cutoff parameter of the viscosity.
@@ 5 Discussion and Conclusion
In this paper we commenced the rigorous analytical study of one of the fundamental models of sea-ice dynamics, namely the elastic-viscous-plastic (EVP) sea-ice model. We introduced a regularised version of the EVP model, by means of the Voigt regularisation, which is inspired by the original elastic regularisation of the Hibler model in Ref. [39] (as both regularisations formally do not impact the asymptotic dynamics). We have shown this Voigt-EVP model to be globally well-posed. One interesting question, which we leave to future work, is whether the Voigt-EVP model also has a suitable notion of weak solution.
Due to the (formal) linear ill-posedness of the EVP model as discussed in Section 1.2, it seems necessary to introduce a regularised version of the EVP model of the type (12a)-(12c), in order to obtain a well-posed sea-ice model. Finally, in order to conclude the proof of the global well-posedness we took the limit of the viscosity cutoff parameter $\epsilon$ going to zero (while keeping the Voigt regularisation), and obtain that the Voigt-EVP model is globally well-posed.
Moreover, we would like to mention that over time several related models to the EVP model have been developed in the sea-ice literature, see for example Refs. [7, 38]. The approach developed here will also have potential applications in the rigorous study of several of these related sea-ice models.
@@
If one includes an advection term in the momentum equation of the intermediate system (13a)-(13c), by applying the same methods as in this paper, one can prove the local well-posedness of the intermediate system by using similar estimates as in this paper. Note however that the existence time will most likely depend on $\epsilon$ (unlike for Corollary 1.4). This therefore prohibits taking the limit $\epsilon\rightarrow 0$, as we have done in the proof of Theorem 1.2.
@@
We remark that the replacement of $\overline{\mathcal{D}}$ by $\lvert D(u)\rvert$ in the Voigt-EVP model is done merely for the sake of convenience. In fact, by using the approach from this paper one can prove the global well-posedness of the model (13a)-(13c) in which $\mathcal{D}_{\epsilon}$ has been replaced by
@@
$\left\lVert\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\lesssim\lVert\nabla u\rVert_{L^{2}}+\epsilon,\quad\left\lVert\nabla\left(\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right)\right\rVert_{L^{2}}\lesssim\lVert\Delta u\rVert_{L^{2}}.$ (79)
@@
We also note that, in addition to (14), different regularisations of the strain rate have been used in the sea-ice literature. For example, one can set maximal values for the shear and bulk viscosities, as is done in Ref. [36]. This means that, instead of $\mathcal{D}_{\epsilon}$, one uses
$\max\{\overline{\mathcal{D}},\epsilon\}=\frac{\overline{\mathcal{D}}+\epsilon}{2}+\frac{1}{2}\lvert\overline{\mathcal{D}}-\epsilon\rvert$. (80)
@@
$\overline{\mathcal{D}}_{\epsilon,\gamma}\coloneqq\frac{\overline{\mathcal{D}}+\epsilon+\gamma}{2}+\frac{1}{2}\sqrt{\lvert\overline{\mathcal{D}}-\epsilon\rvert^{2}+\gamma^{2}}.$ (81)
@@
It is straightforward to repeat the proof of Theorem 1.2 to obtain a global strong solution of the Voigt-EVP model with the viscosity regularisation from Eq. (80) by taking the limit $\gamma\rightarrow 0$. Moreover, one can also choose to regularise the viscosities by means of the hyperbolic tangent, as was introduced in Ref. [52], see also Ref. [18], but again one can treat this regularisation by similar methods as in this paper. Generally speaking, one can treat any $C^{2}$ approximation of $\overline{\mathcal{D}}$ by relying on the approach of this paper.

### 20 Mar 2026 v3

@@ 3.2 Uniform-in-$(N,\epsilon)$ estimates
$\mathcal{T}_{w,i}$ $=c_{w}\rho_{w}\lvert U_{w}-u_{i}\rvert\big[(U_{w}-u_{i})\cos\theta+(U_{w}-u_{i})^{\perp}\sin\theta\big],$ (4.8)
$\mathcal{D}_{i}$ $=\lvert D(u_{i})\rvert$. (4.9)
Subsequently, by proceeding entirely analogously to section 3.3 (the only difference being that we use Lemma 2.3 in the case $\epsilon=0$ rather than $\epsilon>0$), we can obtain the analogous version of estimate (3.44), which is given by
$\begin{gathered}\frac{1}{2}\frac{\rm d}{\ {\rm d}t}\bigg[\lVert\delta u\rVert_{L^{2}}^{2}+\lVert\nabla\delta u\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\lVert\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}(1+\alpha^{2})\lVert\nabla\delta\sigma\rVert_{L^{2}}^{2}+\mathcal{E}^{-1}\alpha^{2}\lVert\Delta\delta\sigma\rVert_{L^{2}}^{2}\bigg]\\ \lesssim(1+\|\sigma_{1}\|_{H^{2}}+\|\sigma_{2}\|_{H^{2}}+\|u_{1}\|_{H^{2}}+\|u_{2}\|_{H^{2}})(\|\delta u\|_{H^{1}}^{2}+\|\delta\sigma\|_{H^{2}}^{2}).\end{gathered}$ (4.10)
@@
We remark that the proof of the convergence of the solutions in the viscosity cutoff limit $\epsilon\rightarrow 0$, is in some sense related to the approach in [3], in which the local well-posedness of a model for non-Newtonian incompressible flows with power-law stress tensor was established uniformly in the cutoff parameter of the viscosity.
@@ 5 Discussion and conclusion
In this paper we commenced the rigorous analytical study of one of the fundamental models of sea-ice dynamics, namely the elastic-viscous-plastic (EVP) sea-ice model. We introduced a regularised version of the EVP model, by means of the Voigt regularisation, which is inspired by the original elastic regularisation of the Hibler model in [38] (as both regularisations formally do not impact the asymptotic dynamics). We have shown this Voigt-EVP model to be globally well-posed. One interesting question, which we leave to future work, is whether the Voigt-EVP model also has a suitable notion of weak solution.
Due to the (formal) linear ill-posedness of the EVP model as discussed in Section 1.2, it seems necessary to introduce a regularised version of the EVP model of the type (1.12a)-(1.12c), in order to obtain a well-posed sea-ice model. Finally, in order to conclude the proof of the global well-posedness we took the limit of the viscosity cutoff parameter $\epsilon$ going to zero (while keeping the Voigt regularisation), and obtain that the Voigt-EVP model is globally well-posed.
Moreover, we would like to mention that over time several related models to the EVP model have been developed in the sea-ice literature, see for example [40, 7]. The approach developed here will also have potential applications in the rigorous study of several of these related sea-ice models.
@@
If one includes an advection term in the momentum equation of the intermediate system (1.13a)-(1.13c), by applying the same methods as in this paper, one can prove the local well-posedness of the intermediate system by using similar estimates as in this paper. Note however that the existence time will most likely depend on $\epsilon$ (unlike for Corollary 1.4). This therefore prohibits taking the limit $\epsilon\rightarrow 0$, as we have done in the proof of Theorem 1.2.
@@
We remark that the replacement of $\overline{\mathcal{D}}$ by $\lvert D(u)\rvert$ in the Voigt-EVP model is done merely for the sake of convenience. In fact, by using the approach from this paper one can prove the global well-posedness of the model (1.13a)-(1.13c) in which $\mathcal{D}_{\epsilon}$ has been replaced by
@@
$\left\lVert\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right\rVert_{L^{2}}\lesssim\lVert\nabla u\rVert_{L^{2}}+\epsilon,\quad\left\lVert\nabla\left(\sqrt{\overline{\mathcal{D}}^{2}+\epsilon^{2}}\right)\right\rVert_{L^{2}}\lesssim\lVert\Delta u\rVert_{L^{2}}.$ (5.1)
@@
We also note that, in addition to (1.14), different regularisations of the strain rate have been used in the sea-ice literature. For example, one can set maximal values for the shear and bulk viscosities, as is done in [36]. This means that, instead of $\mathcal{D}_{\epsilon}$, one uses
$\max\{\overline{\mathcal{D}},\epsilon\}=\frac{\overline{\mathcal{D}}+\epsilon}{2}+\frac{1}{2}\lvert\overline{\mathcal{D}}-\epsilon\rvert$. (5.2)
@@
$\overline{\mathcal{D}}_{\epsilon,\gamma}\coloneqq\frac{\overline{\mathcal{D}}+\epsilon+\gamma}{2}+\frac{1}{2}\sqrt{\lvert\overline{\mathcal{D}}-\epsilon\rvert^{2}+\gamma^{2}}.$ (5.3)
@@
It is straightforward to repeat the proof of Theorem 1.2 to obtain a global strong solution of the Voigt-EVP model with the viscosity regularisation from equation (5.2) by taking the limit $\gamma\rightarrow 0$. Moreover, one can also choose to regularise the viscosities by means of the hyperbolic tangent, as was introduced in [52], see also [17], but again one can treat this regularisation by similar methods as in this paper. Generally speaking, one can treat any $C^{2}$ approximation of $\overline{\mathcal{D}}$ by relying on the approach of this paper.